# Parallel Test-Time Scaling for Latent Reasoning Models

### v1

@@ Abstract
Parallel test-time scaling (TTS) is a pivotal approach for enhancing large language models (LLMs), typically by sampling multiple token-based chains-of-thought in parallel and aggregating outcomes through voting or search. Recent advances in latent reasoning, where intermediate reasoning unfolds in continuous vector spaces, offer a more efficient alternative to explicit Chain-of-Thought, yet whether such latent models can similarly benefit from parallel TTS remains open, mainly due to the absence of sampling mechanisms in continuous space, and the lack of probabilistic signals for advanced trajectory aggregation. This work enables parallel TTS for latent reasoning models by addressing the above issues. For sampling, we introduce two uncertainty-inspired stochastic strategies: Monte Carlo Dropout and Additive Gaussian Noise. For aggregation, we design a Latent Reward Model (LatentRM) trained with step-wise contrastive objective to score and guide latent reasoning. Extensive experiments and visualization analyses show that both sampling strategies scale effectively with compute and exhibit distinct exploration dynamics, while LatentRM enables effective trajectory selection. Together, our explorations open a new direction for scalable inference in continuous spaces. Code released at here.
Parallel Test-Time Scaling for Latent Reasoning Models
Runyang You1, Yongqi Li1‡, Meng Liu2, Wenjie Wang3, Liqiang Nie4, Wenjie Li1 1The Hong Kong Polytechnic University 2Shandong Jianzhu University 3University of Science and Technology of China 4Harbin Institute of Technology (Shenzhen) runyang.y@outlook.com liyongqi0@gmail.com
Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance on challenging tasks through test-time scaling (TTS), where more inference compute leads to better results. Such scaling is often realized through explicit Chain-of-Thought (CoT) reasoning Wei et al. (2022), where models verbalize intermediate solving steps in natural language, generate long token sequences over which compute can be scaled. One promising direction is to scale in parallel Fu et al. (2025); Qu et al. (2025); Snell et al. (2024), which samples multiple reasoning paths and aggregates outcomes via methods such as majority voting Wang et al. (2022), best-of-$N$  Zhou et al. (2022), or guided search Uesato et al. (2022). Through parallel scaling, models transform extra inference compute directly into stronger capabilities, without any need for parameter updates or retraining.
@@
Moving beyond explicit token-by-token reasoning, recent works show that reasoning can instead unfold in latent space, with vector representations replacing tokens as the intermediate steps in autoregressive generation Li et al. (2025); Chen et al. (2025a). This paradigm, also known as continuous CoT (CCOT), has the potential to match or even surpass explicit CoT while being more compact and efficient, resembling intuitive human reasoning Hao et al. (2024); Shen et al. (2025); Tan et al. (2025); Wu et al. (2025). By compressing or distilling explicit reasoning into continuous representations, these methods not only speed up inference but also capture abstract patterns difficult to express in language, positioning latent reasoning as a promising direction for advancing LLM capabilities.
@@
Given the promising potential of latent reasoning, and that parallel TTS can effectively utilize additional generated tokens to deliver superior performance Wang et al. (2022); Muennighoff et al. (2025); Snell et al. (2024), a natural question arises: can latent reasoning models also benefit from parallel TTS? Extending parallel TTS into latent reasoning models is appealing but non-trivial. First, latent reasoning models lack the fundamental sampling capability: while token-based models generate logits that enable commonly-used sampling strategies like top-k or nucleus sampling–as illustrated in Figure 1(a), latent models operate on continuous vectors without an explicit probability distribution, making it difficult to produce diverse reasoning paths. Second, the aggregation mechanism used in token-based models does not directly apply to latent reasoning models: while token-based methods typically utilize token-level probabilities to rank reasoning trajectories, latent reasoning models provide no inherent likelihoods or stepwise scores, making it difficult to evaluate or aggregate the sampled latent reasoning paths.
@@
To address these challenges, we re-think both sampling and aggregation for latent reasoning, proposing effective solutions tailored to the continuous setting. For sampling, to ensure informativeness and controllability of the sampling space, we draw on uncertainty estimation theory Gawlikowski et al. (2023); Liu et al. (2019) and propose two simple yet effective strategies to introduce stochasticity into latent reasoning: Monte Carlo dropout (MC-dropout) to capture epistemic uncertainty; and Additive Gaussian Noise (AGN) to simulate aleatoric uncertainty, as illustrated in Figure 1. For aggregation, we propose the Latent Reward Model (LatentRM), a dedicated scorer that evaluates and guides the progression of latent reasoning at each intermediate step. LatentRM is trained via a step-wise contrastive objective that discriminates among candidate thoughts at each reasoning step, enabling fine-grained, position-sensitive scoring.
@@
(a) Token-based Sampling
(b) Monte Carlo Dropout
(c) Additive Gaussian Noise
Figure 1: Sampling mechanisms for token-based generation (1(a)) and our proposed approaches for the latent setting (1(b) & 1(c)). (1(a)): Multinomial sampling over token probabilities at each step. (1(b)): Monte Carlo Dropout (MC-dropout): Stochastic inference via randomly sampled dropout masks. (1(c)): Additive Gaussian Noise (AGN): independent Gaussian perturbations injected to each latent thought.
@@ 2.1 Test-Time Scaling
Large language models increasingly rely on test-time scaling (TTS)—allocating more computation at inference to improve reasoning quality. Two main axes have emerged. The sequential axis advances by generating longer reasoning traces Muennighoff et al. (2025); Wang et al. (2025), a strategy employed in recent reasoning-oriented models DeepSeek-AI et al. (2024); Yang et al. (2025); You et al. (2025); Fu et al. (2025); Lee et al. (2025). The parallel axis generates multiple reasoning trajectories and aggregates them, as in Self-Consistency Wang et al. (2022), Best-of-N Zhou et al. (2022), and tree-structured exploration methods Uesato et al. (2022). These two axes are often combined, enabling flexible accuracy–compute trade-offs under deployment constraints. Recent works have sought to improve the efficiency of TTS primarily by refining aggregation strategies and decision rules, such as adaptive voting and confidence-based stopping criteria Snell et al. (2024); Brown et al. (2024). Yet all of these approaches ultimately rest on token-level sampling, decoding by drawing tokens from the predictive distribution. Such mechanism is intrinsic to token-based reasoning, but is fundamentally incompatible with latent reasoning, which operates over continuous representations.
@@ 2.2 Latent Reasoning
Chain-of-Thought (CoT) reasoning has proven highly effective, yet it remains constrained by its reliance on natural language Zhu et al. (2025). This introduces inefficiency–most tokens add little value and fail to capture cognition like abstract insights or intuitive leaps, making step-by-step verbalization an unnatural limit. These limitations motivate the shift toward latent reasoning (Chen et al., 2025c, a). One line of work pursues architectural modifications, such as dynamically skipping or repeating transformer layers, or introducing auxiliary modules to adjust computational cost Chen et al. (2025b); Geiping et al. (2025); Mohtashami et al. (2024). More recently, research has shifted to latent autoregressive reasoning, where models build trajectories directly in hidden space. COCONUT Hao et al. (2024) pioneered latent generation using the last hidden state as the next thought, training via curriculum learning, while CODI Shen et al. (2025) introduced latent auto-regression through self-distillation. More recently, CoLaR Tan et al. (2025) applied reinforcement learning with dynamic latent compression. CoLaR Tan et al. (2025) and SoftCoT++ Xu et al. (2025) made initial attempts to inject noise for stochastic latent reasoning, but these efforts remain preliminary. Overall, existing methods leave open the question of how latent reasoning can fully benefit from stochastic exploration and TTS. To address this gap, we develop and systematically analyze stochastic sampling and aggregation strategies in latent space, establishing foundational methods and findings for parallel TTS with latent reasoning models.
@@ 3 Preliminaries
Different latent reasoning models control the reasoning length in distinct ways. COCONUT Hao et al. (2024) and CODI Shen et al. (2025) use predetermined sequence lengths, where a fixed $T$ is set and the <|EoT|> is manually inserted. CoLaR Tan et al. (2025) introduces dynamic latent compression: the reasoning length (or “thinking speed”) is adjusted by a compression factor specified in the prompt (e.g., “thinking speed = 5×”), allowing the model to adaptively decide when to emit the EoT token and terminate reasoning.
@@ 4.1 Stochastic Sampling in Latent Space
We now detail the sampling procedures, and their grounding in uncertainty, with formal derivations in Appendix A.
@@
At inference, we keep dropout active with rate $p$ by sampling a binary mask $m^{(n)}\sim\text{Bernoulli}(p)$ and applying it to the model weights $\bm{\theta}^{(n)}$ during each forward pass, as illustrated in Figure 1(b). Formally, this yields
@@
Following prior practice (Gao et al., 2021; Vaswani et al., 2017), dropout is applied after the feed-forward layer within each Transformer block. Such placement captures epistemic uncertainty, which is the variability arising from the model itself, as each thought corresponds to a different plausible configuration of its weights.
@@
This procedure introduces variance controlled solely by the noise scale $\sigma$, independent of the model parameters, thereby modeling aleatoric uncertainty.
@@ 4.2 Latent Trajectories Aggregation
A key challenge in extending parallel TTS to latent reasoning lies in aggregation. Unlike token-based TTS, where log-likelihoods provide a natural scoring mechanism, latent trajectories are continuous vectors without explicit scores. This prevents direct application of best-of-$N$ or beam search. While process reward models (PRMs) Zhang et al. (2024); Wang et al. (2024) could be a natural choice for evaluating intermediate reasoning steps, latent thoughts are abstract vectors without linguistic form and cannot be interpreted or scored by existing PRMs. To this end, we propose the Latent Reward Model (LatentRM), a dedicated reward model for latent thoughts that enables effective aggregation strategies.
@@
We extend the latent reasoning backbone with an additional scoring head that maps hidden states to a scalar. This scoring head, LatentRM $g_{\bm{\phi}}$, takes in the input prompt $x$ and the generated latent trajectory up to step $t$, and outputs a score:
@@
During inference, LatentRM evaluates each candidate trajectory by computing the sum of logits $\sum_{t}r_{t}$ over the generated sequence, serving as a proxy for the relative quality of thought $\bm{h}_{1:t}$. This logit-summing strategy is justified in Appendix B, which shows that cumulative logits can solely determine trajectory ranking.
@@ 5 Experiment Setup
Following previous works Shen et al. (2025); Tan et al. (2025); Hao et al. (2024), evaluation is conducted on three benchmarks: (1) GSM8K-Test, the official test split of GSM8K comprising 1.3k samples; (2) GSM8K-Hard Gao et al. (2022), a more challenging version of GSM8K-Test where numbers are scaled to larger magnitudes to increase problem difficulty, and (3) MultiArith Uesato et al. (2022), a dataset with 600 test samples focusing on multi-step arithmetic reasoning.
@@
We evaluate three representative latent reasoning models: (1) COCONUT (Hao et al., 2024), which progressively replaces CoT steps with latent thoughts; (2) CODI (Shen et al., 2025), which self-distills CoT into latent space; and (3) CoLaR (Tan et al., 2025), which performs dynamic latent compression with reinforcement learning. All experiments leverage officially released checkpoints, where COCONUT and CODI are backboned on GPT-2, and CoLaR on Llama-3.2-1B. Following the default configurations in the original papers, for COCONUT and CODI, we fix the number of latent thoughts to $T=6$; for CoLaR, thinking speed set to $2$ with a maximum upper limit of 64 latent thoughts.
@@ 5.2 Sampling Evaluation
This metric is equivalent to pass@k when sampling size $N$ and cutoff $k$ are identical Zhang et al. (2024); Brown et al. (2024). Higher coverage indicates stronger sampling effectiveness—i.e., a greater ability to uncover correct reasoning paths as the inference budget increases. All sampling schemes are evaluated under equal $N$ for fair comparison.
@@ 5.3 Aggregation Evaluation
We evaluate the proposed LatentRM through two of its supported aggregation strategies: (1) best-of-$N$ selection scored by LatentRM, (2) beam search guided by LatentRM, compared against majority voting as a non-parametric baseline. To ensure fairness, we match all methods under the same compute budget: best-of-$N$ and majority voting use $N$ independent samples, while beam search adopts a beam size of $\sqrt{N}$ for comparable decoding cost Zhang et al. (2024); Snell et al. (2024). Training configurations and optimization details for LatentRM are provided in Appendix D, and full inference procedures for aggregation strategies are described in Appendix C.
@@ 6.1 Main Results
(a) GSM-Test
(b) GSM-Hard
(c) MultiArith
Figure 2: Coverage (%) versus N plot for COCONUT, CODI, and CoLaR on GSM-Test (2(a)), GSM-Hard (2(b)) and MultiArith (2(c)). Each subplot compares MC-dropout and AGN using the optimal hyperparameter. Higher coverage indicates a larger fraction of problems solved by NN attempts. Results are reported as the mean over three runs.
We systematically evaluate both stochastic sampling strategies by varying the sample count (N) and measuring solution coverage. We tune the AGN hyperparameter $\sigma$ over $[0.01,1.5]$ and the MAC-Dropout probability $p$ over $[0,1]$, using binary search within each interval to maximize coverage@$64$. We then plot two curves using the optimal hyperparameters for each method. Figure 2 presents four key findings: (1) both MC-dropout and AGN can effectively scale, as coverage increases monotonically with larger sample sizes; (2) the marginal improvement diminishes as $N$ grows, suggesting a saturation effect where additional samples contribute less to coverage gains; (3) increased sampling narrows inter-model performance gaps. Notably, at N=64, COCONUT and CODI achieve nearly equivalent coverage, despite CODI’s clear superiority at N=1. (4) MC-dropout consistently achieves higher coverage across nearly all $N$, highlighting its advantage as a more reliable stochastic sampling approach.

### v3

@@ Abstract
Parallel test-time scaling (TTS) is a pivotal approach for enhancing large language models (LLMs), typically by sampling multiple token-based chains-of-thought in parallel and aggregating outcomes through voting or search. Recent advances in latent reasoning, where intermediate reasoning unfolds in continuous vector spaces, offer a more efficient alternative to explicit Chain-of-Thought, yet whether such latent models can similarly benefit from parallel TTS remains open, mainly due to the absence of sampling mechanisms in continuous space, and the lack of probabilistic signals for advanced trajectory aggregation. This work enables parallel TTS for latent reasoning models by addressing the above issues. For sampling, we introduce two uncertainty-inspired stochastic strategies: Monte Carlo Dropout and Additive Gaussian Noise. For aggregation, we design a Latent Reward Model (LatentRM) trained with step-wise contrastive objective to score and guide latent reasoning. Extensive experiments and visualization analyses show that both sampling strategies scale effectively with compute and exhibit distinct exploration dynamics, while LatentRM enables effective trajectory selection. Together, our explorations open a new direction for scalable inference in continuous spaces.
{teaserfigure}
Sampling mechanisms for token-based generation (1(a)) and our proposed approaches for the latent setting (1(b) & 1(c)). (1(a)): Multinomial sampling over token probabilities. (1(b)): Monte Carlo Dropout (MC-dropout): Inference via randomly sampled dropout masks. (1(c)): Additive Gaussian Noise (AGN): independent Gaussian perturbations.
(a) Token-based Sampling
(b) Monte Carlo Dropout
(c) Additive Gaussian Noise
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance on challenging tasks through test-time scaling (TTS), where more inference compute leads to better results. Such scaling is often realized through explicit Chain-of-Thought (CoT) reasoning [wei_chain--thought_2022], where models verbalize intermediate solving steps in natural language, generate long token sequences over which compute can be scaled. One promising direction is to scale in parallel [fu_deep_2025, qu_optimizing_2025, snell_scaling_2024], which samples multiple reasoning paths and aggregates outcomes via methods such as majority voting [wang_self-consistency_2022], best-of-$N$ [zhou_least--most_2022], or guided search [uesato_solving_2022]. Through parallel scaling, models transform extra inference compute directly into stronger capabilities, without any need for parameter updates or retraining.
@@
Moving beyond explicit token-by-token reasoning, recent works show that reasoning can instead unfold in latent space, with vector representations replacing tokens as the intermediate steps in autoregressive generation [li_implicit_2025, chen_reasoning_2025]. This paradigm, also known as continuous CoT (CCOT), has the potential to match or even surpass explicit CoT while being more compact and computationally efficient, resembling intuitive human reasoning [COCONUT, shen_codi_2025, COLAR, wu_parallel_2025]. By effectively compressing or distilling explicit reasoning into continuous representations, these methods not only speed up inference but also capture abstract patterns difficult to express in natural language, positioning latent reasoning as a promising direction for advancing LLM capabilities.
@@
Given the promising potential of latent reasoning, and that parallel TTS can effectively utilize additional generated tokens to deliver superior performance [wang_self-consistency_2022, muennighoffS1SimpleTesttime2025, snell_scaling_2024], a natural question arises: can latent reasoning models also benefit from parallel TTS? Extending parallel TTS into latent reasoning models is appealing but non-trivial. First, latent reasoning models lack the fundamental sampling capability: while token-based models generate logits that enable commonly-used sampling strategies like top-k or nucleus sampling–as illustrated in Figure 1(a), latent models operate on continuous vectors without an explicit probability distribution and therefore .
Second, the aggregation mechanism used in token-based models does not directly apply to latent reasoning models: while token-based methods typically utilize token-level probabilities to rank reasoning trajectories, latent reasoning models provide no inherent likelihoods or stepwise scores, making it difficult to evaluate or aggregate the sampled latent reasoning paths.
@@
To address these challenges, we re-think both sampling and aggregation for latent reasoning, proposing effective solutions tailored to the continuous setting. For sampling, to ensure informativeness and controllability of the sampling space, we draw on uncertainty estimation theory [gawlikowski_survey_2023, NEURIPS2019_1cc8a8ea] and propose two simple yet effective strategies to introduce stochasticity into latent reasoning: Monte Carlo dropout (MC-dropout) to capture epistemic uncertainty; and Additive Gaussian Noise (AGN) to simulate aleatoric uncertainty, as illustrated in Figure Parallel Test-Time Scaling for Latent Reasoning Models. For aggregation, we propose the Latent Reward Model (LatentRM), a dedicated scorer that evaluates and guides the progression of latent reasoning at each intermediate step. LatentRM is trained via a step-wise contrastive objective that discriminates among candidate thoughts at each reasoning step, enabling fine-grained, position-sensitive scoring.
@@ 2.1 Test-Time Scaling
Large language models increasingly rely on test-time scaling (TTS)—allocating more computation at inference to improve reasoning quality. Two main axes have emerged. The sequential axis advances by generating longer reasoning traces [muennighoffS1SimpleTesttime2025, wang_scaling_2025], a strategy employed in recent reasoning-oriented models [deepseek-ai_deepseek-v3_2024, yang_qwen3_2025, you_textr2textec_2025, fu_deep_2025, lee_evolving_2025]. The parallel axis generates multiple reasoning trajectories and aggregates them, as in Self-Consistency [wang_self-consistency_2022], Best-of-N [zhou_least--most_2022], and tree-structured exploration methods [uesato_solving_2022]. These two axes are often combined, enabling flexible accuracy–compute trade-offs under deployment constraints. Recent works have sought to improve the efficiency of TTS primarily by refining aggregation strategies and decision rules, such as adaptive voting and confidence-based stopping criteria [snell_scaling_2024, brown_large_2024]. Yet all of these approaches ultimately rest on token-level sampling, decoding by drawing tokens from the predictive distribution. Such mechanism is intrinsic to token-based reasoning, but is fundamentally incompatible with latent reasoning, which operates over continuous representations.
@@ 2.2 Latent Reasoning
Chain-of-Thought (CoT) reasoning has proven highly effective, yet it remains constrained by its reliance on natural language [zhu_survey_2025]. This introduces inefficiency–most tokens add little value and fail to capture cognition like abstract insights or intuitive leaps, making step-by-step verbalization an unnatural limit. These limitations motivate the shift toward latent reasoning [chen_emergence_2025, chen_reasoning_2025]. One line of work pursues architectural modifications, such as dynamically skipping or repeating transformer layers, or introducing auxiliary modules to adjust computational cost [chen_inner_2025, geiping_scaling_2025, mohtashami_cotformer_2024]. More recently, research has shifted to latent autoregressive reasoning, where models build trajectories directly in hidden space. COCONUT [COCONUT] pioneered latent generation using the last hidden state as the next thought, training via curriculum learning, while CODI [shen_codi_2025] introduced latent auto-regression through self-distillation. While CoLaR [COLAR] and SoftCoT++ [Softcot++] made initial attempts to inject noise for stochastic latent reasoning, these efforts remain preliminary.
@@ 3 Preliminaries
Different latent reasoning models control the reasoning length in distinct ways. COCONUT [COCONUT] and CODI [shen_codi_2025] use predetermined sequence lengths, where a fixed $T$ is set and the <|EoT|> is manually inserted. CoLaR [COLAR] introduces dynamic latent compression: the reasoning length (or “thinking speed”) is adjusted by a compression factor specified in the prompt (e.g., “thinking speed = 5×”), allowing the model to adaptively decide when to emit the EoT token and terminate reasoning.
@@ 4.1 Stochastic Sampling in Latent Space
We now detail the sampling procedures. Formal derivations and algorithmic details are provided in Appendix B and Appendix H, respectively.
@@
We keep dropout active at inference with rate $p$, sampling binary masks $m^{(n)}\sim\text{Bernoulli}(p)$ applied to model weights $\bm{\theta}^{(n)}$:
@@
Dropout is applied after the feed-forward layer in each Transformer block [gao_simcse_2021], capturing epistemic uncertainty as each pass samples a different weight configuration.
@@
This procedure models aleatoric uncertainty, as the variance is controlled solely by the noise scale $\sigma$, independent of the model parameters.
@@ 4.2 Latent Trajectories Aggregation
A key challenge in extending parallel TTS to latent reasoning lies in aggregation. Unlike token-based TTS, where log-likelihoods provide a natural scoring mechanism, latent trajectories are continuous vectors without explicit scores. This prevents direct application of best-of-$N$ or beam search. While process reward models (PRMs) [zhang_openprm_2024, math-shepherd] could be a natural choice for evaluating intermediate reasoning steps, latent thoughts are abstract vectors without linguistic form and cannot be interpreted or scored by existing PRMs. To this end, we propose the Latent Reward Model (LatentRM), a dedicated reward model for latent thoughts that enables effective aggregation strategies.
@@
We extend the latent reasoning backbone with an additional scoring head that maps hidden states to a scalar. LatentRM $g_{\bm{\phi}}$ takes in the input prompt $x$ and the generated latent trajectory up to step $t$ and outputs a score:
@@
During inference, LatentRM evaluates each candidate trajectory by computing the sum of logits $\sum_{t}r_{t}$ over the generated sequence, serving as a proxy for the relative quality of thought $\bm{h}_{1:t}$. This logit-summing strategy is justified in Appendix C, which shows that cumulative logits can solely determine trajectory ranking.
@@ 5 Experiment Setup
Following previous works [shen_codi_2025, COLAR, COCONUT], evaluation is conducted on three benchmarks: (1) GSM8K-Test, the official test split of GSM8K (1,319 test samples) [^1] https://huggingface.co/datasets/gsm8k; (2) GSM8K-Hard [GSM-HARD], a more challenging version of GSM8K-Test where numbers are scaled to larger magnitudes to increase problem difficulty with (1,319 test samples) [^2] https://huggingface.co/datasets/reasoning-machines/gsm-hard, and (3) MultiArith [uesato_solving_2022], a dataset focusing on multi-step arithmetic reasoning (600 test samples) [^3] https://huggingface.co/datasets/lighteval/multi_arith.
@@
We evaluate three representative latent reasoning models: (1) COCONUT [COCONUT], which progressively replaces CoT steps with latent thoughts; (2) CODI [shen_codi_2025], which self-distills CoT into latent space; and (3) CoLaR [COLAR], which performs dynamic latent compression with reinforcement learning. All experiments leverage officially released checkpoints, where COCONUT and CODI are backboned on GPT-2 , and CoLaR on Llama-3.2-1B. Following the default configurations in the original papers, for COCONUT and CODI, we fix the number of latent thoughts to $T=6$; for CoLaR, thinking speed set to $2$ with a maximum upper limit of 64 latent thoughts.
@@ 5.2 Sampling Evaluation
This metric is equivalent to pass@k when sampling size $N$ and cutoff $k$ are identical [zhang_openprm_2024, brown_large_2024]. Higher coverage indicates stronger sampling effectiveness—i.e., a greater ability to uncover correct reasoning paths as the inference budget increases. All sampling schemes are evaluated under equal $N$ for fair comparison.
@@ 5.3 Aggregation Evaluation
We evaluate the proposed LatentRM through two of its supported aggregation strategies: (1) best-of-$N$ selection scored by LatentRM, (2) beam search guided by LatentRM, compared against majority voting as a non-parametric baseline. To ensure fairness, we match all methods under the same compute budget: best-of-$N$ and majority voting use $N$ independent samples, while beam search adopts a beam size of $\sqrt{N}$ for comparable decoding cost [zhang_openprm_2024, snell_scaling_2024]. Training configurations and optimization details for LatentRM are provided in Appendix G, and full inference procedures for aggregation strategies are described in Appendix D.
(d) GSM-Test
(e) GSM-Hard
(f) MultiArith
Figure 1:
@@ 6.1 Main Results
We systematically evaluate both stochastic sampling strategies by varying the sample count (N) and measuring solution coverage. We tune the AGN hyperparameter $\sigma$ over $[0.01,1.5]$ and the MAC-Dropout probability $p$ over $[0,1]$, using binary search within each interval to maximize coverage@$64$. We then plot two curves using the optimal hyperparameters for each method. Additonal results on harder benchmarks are provided in Appendix E. Figure 1 presents four key findings: (1) both MC-dropout and AGN can effectively scale, as coverage increases monotonically with larger sample sizes; (2) the marginal improvement diminishes as $N$ grows, suggesting a saturation effect where additional samples contribute less to coverage gains; (3) increased sampling narrows inter-model performance gaps. Notably, at N=64, COCONUT and CODI achieve nearly equivalent coverage, despite CODI’s clear superiority at N=1. (4) MC-dropout consistently achieves higher coverage across nearly all $N$, highlighting its advantage as a more reliable stochastic sampling approach.